# Invariant measures for some dissipative systems from the Jacobi last multiplier

### v1

@@
Hamiltonian dynamics describing conservative systems naturally preserves the standard notion of phase-space volume, a result known as the Liouville’s theorem which is central to the formulation of classical statistical mechanics. In this paper, we obtain explicit expressions for invariant phase-space measures for certain dissipative systems, namely, systems described by conformal vector fields on symplectic manifolds that are cotangent bundles, contact Hamiltonian systems, and systems of the Liénard class with position-dependent damping. The latter class of systems can be described by certain generalized conformal vector fields on the cotangent bundle of the configuration space. The computation of the invariant measures is achieved by calculating the Jacobi last multiplier for the above-mentioned dissipative systems.
@@ 1 Introduction
Apart from linearly-damped systems with constant damping strength, another class of systems where the second-order equation of motion breaks time-reversal invariance are systems of the Liénard class, being described by the equation
@@
where $f(x)$ and $g(x)$ are suitable (usually smooth) real-valued functions of the real variable $x$. Here, the damping strength may depend on the variable $x$, thereby giving rise to the possibility of nonlinear dynamics as well as that of limit cycles [14]. The phase-space flows admit a non-trivial divergence unlike conservative Hamiltonian systems respecting the Liouville’s theorem. As may be expected, such systems cannot be described using the standard Hamiltonian approach although some such systems may be described by using nonstandard forms of Lagrangian and Hamiltonian functions [8, 10].
The aim of this paper is to present expressions for invariant phase-space measures for some dissipative systems, focusing on three classes of systems – (a) systems that are described by conformal Hamiltonian dynamics, (b) systems that are described by contact Hamiltonian dynamics, and (c) systems of the Liénard class as defined above. For this purpose, we shall make use of the notion of the Jacobi last multiplier [15]. In particular, we will re-derive the results of Bravetti and Tapias for contact Hamiltonian systems [11] by using the formalism of the Jacobi last multiplier. Further, we shall describe a certain generalization of conformal Hamiltonian dynamics on ‘exact’ symplectic manifolds (specifically, cotangent bundles) which will allow us to discuss Liénard-type systems with position-dependent damping. In particular, it will be shown that when $f(x)$ and $g(x)$ satisfy the so-called Cheillini integrability condition (see for example, [8]), it is possible to find analytical expressions for phase-space measures which are invariant under the corresponding dissipative dynamics.
The paper is organized as follows. In the next section [Sec. (2)], we will recall some basic definitions regarding the Jacobi last multiplier and the Hamiltonian formalism on symplectic and contact manifolds. Following this, in Sec. (3), we shall present invariant phase-space measures for both conformal Hamiltonian dynamics as well as contact Hamiltonian dynamics. Finally, in Sec. (4), we will present a generalized version of conformal Hamiltonian dynamics which can describe Liénard-type systems [Eq. (1.1)] and will also present the explicit expression for invariant phase-space measures when the so-called Cheillini integrability condition is satisfied. We will conclude the paper by summarizing the results in Sec. (5).
@@ 2 Preliminaries
In this section, we will review some well-known notions which will be useful for our analysis in the subsequent sections. This will also help us set the notation and clarify the basic conventions. We will begin with the Jacobi last multiplier below which will be followed by a review of the Hamiltonian frameworks on symplectic and contact manifolds.
@@ 2.1 Jacobi last multiplier
The last multiplier which was introduced by Jacobi in 1844 (see the classic text [15]) has turned out to be a useful tool in analytical mechanics. On one hand, the Jacobi last multiplier allows one to deduce a conserved quantity for a planar dynamical system or equivalently, for a dynamical system on an $m$-dimensional phase space if $(m-2)$ conserved quantities are already known – in that case, the system can be reduced to a planar one [15, 16]. On the other hand, the last multiplier allows one to find the Lagrangians describing certain second-order differential equations [17, 18, 19, 20, 21, 22] (see also, the older work [23]), most notably, some from the Liénard class of systems [21, 24, 25]. More recently, it has found use in the context of non-divergence-free vector fields [26], and also in the context of Lie symmetries [27, 28] and integrability [29].
Consider a dynamical system, i.e., a system of first-order equations which go as $\dot{x}_{i}=X_{i}(x_{i})$, where $i=1,2,\cdots,m$. Here, $x_{i}$ could be thought of as being the (possibly local) coordinates in some region $U\subseteq\mathbb{R}^{m}$ and ${X_{i}}$ are real-valued and differentiable functions defined on $U$. The dynamical system is described by a first-order vector field that goes as
@@
For any function $F$ defined on $U$, we have $X(F)=\frac{dF}{dt}$. Let us consider the volume-form $\Omega=dx_{1}\wedge dx_{2}\wedge\cdots\wedge dx_{m}$. The divergence of $X$ can be defined from the Lie derivative as £$\pounds_{X}\Omega=({\rm div}\cdot X)\Omega$; if the divergence is zero, then the vector field is volume preserving, e.g., symplectic vector fields. In general, ${\rm div}\cdot X\neq 0$. We can now define the Jacobi last multiplier.
@@
The Jacobi last multiplier $M$ is a factor such that $MX$ has zero divergence.
@@
where $X_{i}=X_{i}(x_{i})$.
Proof – Consider the dynamical vector field $X=X_{i}\frac{\partial}{\partial x_{i}}$, where $X_{i}$ are suitable functions of the coordinates $x_{i}$. In local coordinates, the condition ${\rm div}\cdot(MX)=0$ is equivalent to
@@
Proof – It is easy to observe that £$\pounds_{MX}\Omega=0\iff\pounds_{X}(M\Omega)=0$.
@@ 2.2.1 Hamiltonian dynamics
The non-degeneracy of $\omega$ allows one to define a vector-bundle isomorphism between the tangent and cotangent bundles of $\mathcal{M}_{s}$ as
@@
Thus, the integral curves of the vector field $X_{H}$ satisfy the Hamilton’s equations. An interesting consequence of (2.5) is that [^2] Here, one makes use of Cartan’s ‘magic’ formula for the Lie derivative: £$\pounds_{X}\alpha=d(\iota_{X}\alpha)+\iota_{X}(d\alpha),$ (2.9) where $\alpha$ is a differential form and $X$ is a vector field.
@@
Thus, a conservative Hamiltonian system may be formally defined as follows:
@@
A conservative Hamiltonian system is the triple $(\mathcal{M}_{s},\omega,H)$, where $(\mathcal{M}_{s},\omega)$ is a symplectic manifold and $H\in C^{\infty}(\mathcal{M}_{s},\mathbb{R})$. The corresponding dynamics is described by a Hamiltonian vector field which is determined by the condition $\iota_{{X}_{H}}\omega=dH$.
@@ 2.2.2 Conformal Hamiltonian dynamics
Now, for the phase space which is a cotangent bundle with tautological one-form $\theta$, we may write
@@
which, upon plugging into Cartan’s magic formula (2.9) gives (2.11). In the Darboux coordinates (also called canonical coordinates), one has the standard expression of the tautological one-form which reads $\theta=-p_{i}dq^{i}$ and which is consistent with (2.6). This means the vector field ${X}^{\gamma}_{H}$ takes the following appearance:
@@
and which means the equations of motion read
@@
where $X_{H}$ is the standard (conservative) Hamiltonian vector field (2.7) and $\Delta=p_{i}\frac{\partial}{\partial p_{i}}$ is the Liouville vector field. Viewing the phase space as a cotangent bundle $\pi:T^{*}Q\rightarrow Q$ on which $p_{i}$ are the induced fiber coordinates, i.e., $\pi:(q^{i},p_{i})\rightarrow q^{i}$, the vector field $\Delta$ generates dilatations along the fiber. Here, the Darboux coordinates $(q^{i},p_{i})$ on the cotangent bundle $T^{*}Q$ are such that $q^{i}$ are the base-space coordinates (on $Q$) while $p_{i}$ are the fiber coordinates such that $\theta=-p_{i}dq^{i}$. Notice that one has the following relationship between the Liouville vector field and the tautological one-form on a cotangent bundle:
@@ 2.3 Hamiltonian dynamics on contact manifolds
Let us now formally define a contact manifold. A contact manifold is a pair $(\mathcal{M}_{c},\eta)$, where $\mathcal{M}_{c}$ is a smooth manifold of real dimension $12n+$ 1 and $\eta$ is a one-form satisfying
$\eta\wedge(d\eta)^{n}\neq 0,$ (2.17)
where $\eta\wedge(d\eta)^{n}$ is the considered volume-form on $\mathcal{M}_{c}$. In the context of Frobenius integrability, the condition (2.17) means that the hyperplane distribution defined as ${\rm ker}(\eta)$ is maximally non-integrable [^4] A more rigorous definition of a contact manifold relies on the existence of a hyperplane distribution which is maximally non-integrable. Such a distribution can be locally expressed as the kernel of a one-form satisfying the condition (2.17). If one can always write the associated hyperplane distribution as ${\rm ker}(\eta)$ (not just locally), then the contact manifold is called an exact contact manifold. In this paper, we will be considering only exact contact manifolds without further stating it explicitly.. The reader is referred to [1, 6, 7] for more details.
@@
Darboux theorem asserts that if the condition (2.17) is satisfied, then it is possible to define local (Darboux) coordinates $(s,q^{i},p_{i})$ near a point such that
@@
Thus, given some ‘contact Hamiltonian’ function $h\in C^{\infty}(\mathcal{M}_{c},\mathbb{R})$, the relations (2.20) define a map $h\mapsto X_{h}$ by which one can associate with it a contact vector field which yields certain equations of motion. Therefore, we shall define a contact Hamiltonian system as follows:
@@
£$\pounds_{X_{h}}(\eta\wedge(d\eta)^{n})={\rm div}.X_{h}(\eta\wedge(d\eta)^{n}),$ (2.24)
@@ 2.4 Constant-damping systems
$\ddot{q}+\gamma\dot{q}+V^{\prime}(q)=0,\quad\gamma>0,\quad V(q)\in C^{\infty}(\mathbb{R},\mathbb{R}),$ (2.26)
where the prime denotes derivation with respect to $q$. The dynamics can be described by either a conformal vector field on a two-dimensional phase space with a symplectic structure or by a contact vector field on a three-dimensional (enlarged) phase space with a contact structure. Let us consider the two cases one after the other.
@@ 2.4.2 Description via contact vector field
[…]
@@ 3 Invariant measures from Jacobi last multiplier
In this section, we will compute the Jacobi last multiplier leading to the computation of invariant measures on the phase space for conformal Hamiltonian systems as well as for contact Hamiltonian systems. Let us take the two cases one after the other.
@@ 3.1 Conformal Hamiltonian dynamics
$\frac{dH}{dt}=-\gamma\Delta(H),$ (3.1)
and that ${\rm div}\cdot X^{\gamma}_{H}=-\gamma n$. From (2.2), the Jacobi last multiplier should satisfy the following equation:
$\frac{d}{dt}\ln M=\gamma n.$ (3.2)
@@
Consider a conformal Hamiltonian system $(\mathcal{M}_{s},\omega,H,\gamma)$ with ${\rm dim~{}}\mathcal{M}_{s}=2n$. In the region of the phase space where $\Delta(H)\neq 0$, the Jacobi last multiplier reads
@@
$\Omega|_{H\neq 0}=\exp\bigg{(}-\frac{1}{2}\frac{dH}{H}\bigg{)}\omega=\frac{\omega}{H^{1/2}}.$ (3.7)

### v2

@@
Hamiltonian dynamics describing conservative systems naturally preserves the standard notion of phase-space volume, a result known as the Liouville’s theorem which is central to the formulation of classical statistical mechanics. In this paper, we obtain explicit expressions for invariant phase-space measures for certain (generally dissipative) mechanical systems, namely, systems described by conformal vector fields on symplectic manifolds that are cotangent bundles, contact Hamiltonian systems, and systems of the Liénard class. The latter class of systems can be described by certain generalized conformal vector fields on the cotangent bundle of the configuration space. The computation of the invariant measures is achieved by making use of the formalism of Jacobi last multipliers.
@@ 1 Introduction
Apart from linearly-damped systems with constant damping strength, another class of systems where the second-order equation of motion breaks time-reversal invariance are systems of the Liénard class, being described by the equation [^2] It may be noted that despite the presence of the damping term ‘$f(x)\dot{x}$’, the dynamics may not always be dissipative. For example, the nonlinear Liénard system $\ddot{x}+kx\dot{x}+\omega^{2}x+\frac{k^{2}}{9}x^{3}=0$ admits oscillatory (non-decaying) solutions [14].
@@
where $f(x)$ and $g(x)$ are suitable (usually smooth) real-valued functions of the real variable $x$. Here, the damping strength may depend on the variable $x$, thereby giving rise to the possibility of nonlinear dynamics as well as that of limit cycles [15]. The phase-space flows admit a non-trivial divergence unlike conservative Hamiltonian systems respecting traditional Liouville’s theorem. As may be expected, such systems cannot be described using the standard Hamiltonian approach although some such systems may be described by using nonstandard forms of Lagrangian and Hamiltonian functions [8, 10] provided $f(x)$ and $g(x)$ satisfy the so-called Cheillini integrability condition:
$\frac{d}{dx}\bigg{(}\frac{g(x)}{f(x)}\bigg{)}+l(l+1)f(x)=0,$ (1.2)
where $l$ is a real number.
The aim of this paper is to present expressions for invariant phase-space measures for some dissipative systems, focusing on three classes of systems – (a) systems that are described by conformal Hamiltonian dynamics, (b) systems that are described by contact Hamiltonian dynamics, and (c) systems of the Liénard class as defined above. For this purpose, we shall make use of the notion of Jacobi last multipliers [16]. In particular, we will re-derive the so-called ‘canonical’ invariant measure for contact Hamiltonian systems [11] by using the formalism of Jacobi last multipliers. Further, we shall describe a certain generalization of conformal Hamiltonian dynamics on ‘exact’ symplectic manifolds (specifically, cotangent bundles) which will allow us to discuss Liénard-type systems. In particular, it will be shown that when $f(x)$ and $g(x)$ satisfy the Cheillini integrability condition (1.2) (see for example, [8]), it is possible to find analytical expressions for phase-space measures which are invariant under the corresponding dynamics.
The paper is organized as follows. In the next section [Sec. (2)], we will recall some basic definitions regarding Jacobi last multipliers and the Hamiltonian formalism on symplectic and contact manifolds. Following this, in Sec. (3), we shall present invariant phase-space measures for both conformal Hamiltonian dynamics as well as contact Hamiltonian dynamics. Finally, in Sec. (4), we will present a generalized version of conformal Hamiltonian dynamics which can describe Liénard-type systems [Eq. (1.1)] and will also present the explicit expressions for invariant phase-space measures when the Cheillini integrability condition (1.2) is satisfied. We will conclude the paper with an extended discussion in Sec. (5).
@@ 2 Preliminaries
In this section, we will review some well-known notions which will be useful for our analysis in the subsequent sections. This will also help us set the notation and clarify the basic conventions. We will begin by reviewing the framework of Jacobi last multipliers below which will be followed by a review of the Hamiltonian frameworks on symplectic and contact manifolds.
@@ 2.1 Jacobi last multiplier
The last multiplier which was introduced by Jacobi in 1844 (see the classic text [16]) has turned out to be a useful tool in analytical mechanics. On one hand, the notion of Jacobi last multipliers allows one to deduce a conserved quantity (at least, locally) for a planar dynamical system or equivalently, for a dynamical system on an $m$-dimensional phase space if $(m-2)$ conserved quantities are already known – in that case, the system can be reduced to a family of planar ones [16, 17]. On the other hand, a last multiplier allows one to find the Lagrangians describing certain second-order differential equations [18, 19, 20, 21, 22, 23] (see also, the older work [24]), most notably, some from the Liénard class of systems [22, 25, 26]. More recently, it has found use in the context of non-holonomic Lagrangian mechanics [27], non-divergence-free vector fields [28], and also in the context of Lie symmetries [29, 30] and integrability [31].
Consider a dynamical system, i.e., a system of first-order equations which go as $\dot{x}_{i}=X_{i}(x_{1},x_{2},\cdots,x_{m})$, where $i=1,2,\cdots,m$. Here, $x_{i}$ could be thought of as being the (possibly local) coordinates in some region $U\subseteq\mathbb{R}^{m}$ and ${X_{i}}$ are real-valued and differentiable functions defined on $U$. The dynamical system is described by a first-order vector field that goes as
@@
Note that we have the definition $X(F)=\frac{dF}{dt}$, for a function $F$ defined on $U$. Let us consider the volume-form $\Omega=dx_{1}\wedge dx_{2}\wedge\cdots\wedge dx_{m}$. The divergence of $X$ can be defined from the Lie derivative as £$\pounds_{X}\Omega=({\rm div}X)\Omega$; if the divergence is zero, then the vector field is volume preserving. In general, ${\rm div}X\neq 0$. We can now define a Jacobi last multiplier (see [27, 28] for discussions on the definition and results presented below).
@@
A Jacobi last multiplier $M$ is a factor such that $MX$ has zero divergence.
@@
where $X_{i}=X_{i}(x_{1},x_{2},\cdots,x_{m})$.
Proof – Consider the dynamical vector field $X=X_{i}\frac{\partial}{\partial x_{i}}$, where $X_{i}$ are suitable functions of the coordinates $x_{i}$ with $i=1,2,\cdots,m$. In local coordinates, the condition ${\rm div}(MX)=0$ is equivalent to
@@
Proof – Since £$\pounds_{MX}\Omega=0$, from Cartan’s formula for the Lie derivative [^3] Here, one makes use of Cartan’s ‘magic’ formula for the Lie derivative: £$\pounds_{X}\alpha=d(\iota_{X}\alpha)+\iota_{X}(d\alpha),$ (2.4) where $\alpha$ is a differential form and $X$ is a vector field., we have $d(\iota_{MX}\Omega)=\pounds_{X}(M\Omega)=0$.
@@ 2.2.1 Hamiltonian dynamics
The non-degeneracy of $\omega$ allows one to define a vector-bundle isomorphism which induces a $C^{\infty}(\mathcal{M}_{s},\mathbb{R})$-linear isomorphism between the set of vector fields and one-forms on $\mathcal{M}_{s}$. In particular, the Hamiltonian vector field $X_{H}$ is defined by
@@
Thus, the integral curves of the vector field $X_{H}$ satisfy the Hamilton’s equations. An interesting consequence of (2.6) is that by using the Cartan’s formula (2.4), one gets
@@
Thus, a (conservative) Hamiltonian system may be formally defined as follows:
@@
A Hamiltonian system is the triple $(\mathcal{M}_{s},\omega,H)$, where $(\mathcal{M}_{s},\omega)$ is a symplectic manifold and $H\in C^{\infty}(\mathcal{M}_{s},\mathbb{R})$. The corresponding dynamics is described by a Hamiltonian vector field which is determined by the condition $\iota_{{X}_{H}}\omega=dH$.
@@ 2.2.2 Conformal Hamiltonian dynamics
Now, for the phase space which is a cotangent bundle with tautological one-form $\theta$, we may define the vector field $X^{\gamma}_{H}$ alternatively as
@@
which, upon plugging into Cartan’s formula (2.4) gives (2.11). In the Darboux coordinates (also called canonical coordinates), one has the standard expression of the tautological one-form which reads $\theta=-p_{i}dq^{i}$ and which is consistent with (2.7). This means the vector field ${X}^{\gamma}_{H}$ takes the following appearance:
@@
and which means the integral curves of $X^{\gamma}_{H}$ are solutions of the equations
@@
where $X_{H}$ is the standard (conservative) Hamiltonian vector field (2.8) and $\Delta=p_{i}\frac{\partial}{\partial p_{i}}$ is the Liouville vector field. Viewing the phase space as a cotangent bundle $\pi:T^{*}Q\rightarrow Q$ on which $p_{i}$ are the induced fiber coordinates, i.e., $\pi:(q^{i},p_{i})\rightarrow q^{i}$, the vector field $\Delta$ generates dilatations along the fiber at each point. Here, the Darboux coordinates $(q^{i},p_{i})$ on the cotangent bundle $T^{*}Q$ are such that $q^{i}$ are the base-space coordinates (on $Q$) while $p_{i}$ are the fiber coordinates such that $\theta=-p_{i}dq^{i}$. Notice that one has the following relationship between the Liouville vector field and the tautological one-form on a cotangent bundle:
@@ 2.3 Hamiltonian dynamics on contact manifolds
Let us now formally define a contact manifold.
Definition 2.5
A contact manifold is a pair $(\mathcal{M}_{c},\eta)$, where $\mathcal{M}_{c}$ is a smooth manifold of (real) dimension $12n+$ 1 and $\eta$ is a one-form satisfying
$\eta\wedge(d\eta)^{n}\neq 0.$ (2.17)
Remarks: Here, $\eta\wedge(d\eta)^{n}$ is the considered volume-form on $\mathcal{M}_{c}$. In the context of Frobenius integrability, the condition (2.17) means that the hyperplane distribution defined as ${\rm ker}(\eta)$ is maximally non-integrable [^5] A more rigorous definition of a contact manifold relies on the existence of a hyperplane distribution which is maximally non-integrable. Such a distribution can be locally expressed as the kernel of a one-form satisfying the condition (2.17). If one can always write the associated hyperplane distribution as ${\rm ker}(\eta)$ (not just locally), then the contact manifold is called an exact contact manifold. In this paper, we will be considering only exact contact manifolds without further stating it explicitly.. The reader is referred to [1, 6, 7] for more details.
@@
Darboux theorem asserts that if the condition (2.17) is satisfied, then it is possible to define local (Darboux) coordinates $(s,q^{i},p_{i})$ with $i\in\{1,2,\cdots,n\}$ near a point such that
@@
Thus, given some ‘contact Hamiltonian’ function $h\in C^{\infty}(\mathcal{M}_{c},\mathbb{R})$, the relations (2.20) define a map $h\mapsto X_{h}$ by which one can associate with it a contact vector field which yields certain equations of motion. We shall define a contact Hamiltonian system as follows:
@@
£$\pounds_{X_{h}}(\eta\wedge(d\eta)^{n})={\rm div}X_{h}(\eta\wedge(d\eta)^{n}),$ (2.24)
@@ 2.4 Constant-damping systems
$m\ddot{q}+m\gamma\dot{q}+V^{\prime}(q)=0,\quad\gamma>0,\quad V(q)\in C^{\infty}(\mathbb{R},\mathbb{R}),$ (2.26)
where the prime denotes derivation with respect to $q$. The dynamics can be described by either a conformal vector field on a two-dimensional phase space $T^{*}\mathbb{R}$ endowed with a symplectic structure or a contact vector field on a three-dimensional (enlarged) phase space $T^{*}\mathbb{R}\times\mathbb{R}$ endowed with a contact structure. Let us consider the two cases one after the other.
@@ 2.4.2 Description via contact vector field
The relation $\dot{q}=p/m$ implies that $p$ is the linear momentum if one interprets $q$ as a mechanical (linear) coordinate. Combining this with the equation of motion for $p$ implies (2.26) which represents the dynamics of a particle moving in a potential but under the influence of linear (constant) damping. If $h$ is independent of $s$, i.e., the Hamiltonian is of the standard type (2.27), we have conservation of $h$ and the preservation of the phase-space volume $\eta\wedge d\eta=ds\wedge dq\wedge dp$.
@@ 3 Invariant measures from Jacobi last multiplier
In this section, we will compute Jacobi last multipliers leading to the computation of invariant measures on the phase space for conformal Hamiltonian systems as well as for contact Hamiltonian systems. Let us take the two cases one after the other.
@@ 3.1 Conformal Hamiltonian dynamics
$X^{\gamma}_{H}(H)=-\gamma\Delta(H),$ (3.1)
and that ${\rm div}X^{\gamma}_{H}=-\gamma n$. From (2.2), a Jacobi last multiplier should satisfy the following equation:
$X^{\gamma}_{H}(\ln M)=\gamma n.$ (3.2)
@@
Consider a conformal Hamiltonian system $(\mathcal{M}_{s},\omega,H,\gamma)$ with ${\rm dim~{}}\mathcal{M}_{s}=2n$. In the region of the phase space where $\Delta(H)\neq 0$, there is a Jacobi last multiplier given by
@@
$\Omega|_{H\neq 0}=\exp\bigg{(}-\frac{1}{2}\int\frac{dH}{H}\bigg{)}\omega=\frac{\omega}{H^{1/2}}.$ (3.7)
Remarks: It turns out that if $V^{\prime}(q)\neq 0$, then the integrand of (3.6) is not an exact differential and consequently, the expression (3.6) is of limited utility except for the free particle. However, for constant-damping systems in a non-trivial potential, invariant measure(s) can be found by considering either a contact-geometric description [Sec (3.2.3) below] or via a description in terms of generalized conformal vector fields as we will show in Sec. (4.3.1); in the latter approach, the potential function must be carefully chosen so as to satisfy the Cheillini integrability condition (1.2).